If both animals are running at full speed,
with the gazelle 55.1 m ahead, how long before
the cheetah catches its prey?
Answer in units of s.
1
Expert's answer
2019-10-14T09:17:29-0400

The speed of cheetah:


"v_c=90.5\frac{km}{h}=\frac{90.5}{3.6}\frac{m}{s}"

The speed of gazelle:


"v_g=74.7\frac{km}{h}=\frac{74.7}{3.6}\frac{m}{s}"

The time will be


"t=\frac{d}{v_c-v_g}=\frac{(55.1)(3.6)}{90.5-74.7}=12.6\ s"